OpenStudy (anonymous):

You construct an open-top box by cutting equal sized squares (x by x inches) out of the four corners of a 14 inch by 24 inch sheet of metal and then folding up the sides. Find a function for the volume of the box as a function of x and then determine the largest volume possible.

OpenStudy (anonymous):

I constructed the box and got v=x * 24x * 14x which came out to v=336x^2 but I dont think I am right. Can anyone please help?

OpenStudy (anonymous):

Sorry I meant v=336^3 but I still dont think its right

OpenStudy (anonymous):

you are stoopid

OpenStudy (anonymous):

Sorry I meant v=336^3 but I still dont think its right